# DHP: Discrete Hierarchical Planning for HRL Agents

### v2

@@ 7. Answer: [Yes]
    Theory assumptions and proofs
  10. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  11. Answer: [Yes]
  12. Justification: All theoretical results include explicit assumptions and complete proofs (Sec. A).
  13. 4.
    Experimental result reproducibility
  14. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  15. Answer: [Yes]
  16. Justification: We include the full training details in the section B that includes all the important hyperparameters, architecture, and optimization objectives for reproducing the results. Although our implementation is based on Director [14], all novel components are self-contained in the algorithms provided (Algs. 1 6). Code release will include a Dockerfile for environment replication and pre-trained checkpoints.
  17. 5.
    Open access to data and code
  18. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  19. Answer: [Yes]
  20. Justification: While we commit to releasing code post-acceptance, the current submission lacks:
    • Anonymized code/scripts in supplemental materials
    • Environment setup instructions (Dockerfile)
    • Exact reproduction commands for key experiments
    Post-release will include:
    • Complete code with environment setup.
    • Pre-trained checkpoints for all tasks.
    • Jupyter notebooks for all results and visualizations.
  21. 6.
    Experimental setting/details
  22. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  23. Answer: [Yes]
  24. Justification: We try to provide all the details specific to our approach in the main paper. We include the remaining details in the supplementary (Sec. B), which include the hyperparameters and the function algorithms implemented.
  25. 7.
    Experiment statistical significance
  26. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  27. Answer: [Yes]
  28. Justification: The statistical descriptions have been added to the paper. For all our plots, the results are mean and standard deviations (std) across $3$ seeds. The bar graphs are shown as mean and std over $100$ evaluation runs. However, we have a small sample size of $3$ seeds due to resource constraints. Error bars represent population standard deviation (not SEM) calculated via numpy.ndarray.std(). We plan $5$-seed runs for experiments for the camera-ready version.
  29. 8.
    Experiments compute resources
  30. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  31. Answer: [Yes]
  32. We include the exact implementation and resource details (time and compute) in the section B.
  33. 9.
    Code of ethics
  34. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  35. Answer: [Yes]
  36. Justification: We do not include any information that can violate anonymity. We try to cite all references to previous work and sources.
  37. 10.
    Broader impacts
  38. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  39. Answer: [Yes]
  40. Justification: We discuss the possible impact of our work in section C. We emphasize that these are speculative, given our simulation-only experiments.
  41. 11.
    Safeguards
  42. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  43. Answer: [N/A]
  44. Justification: We do not introduce any data or models that can be deployed directly.
  45. 12.
    Licenses for existing assets
  46. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  47. Answer: [Yes]
  48. Justification: All reused assets are properly credited with licenses and terms explicitly stated:
    • Director Codebase: Built on the open-source implementation from [14], with modifications documented in the paper (Mentioned in the paper).
    • Miniworld Env: Apache 2.0 license
    No proprietary datasets or models were used.
  49. 13.
    New assets
  50. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  51. Answer: [Yes]
  52. Justification: We will provide the code as an asset with a comprehensive README.
  53. 14.
    Crowdsourcing and research with human subjects
  54. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  55. Answer: [N/A]
  56. Justification: Our work does not involve crowdsourcing nor research with human subjects.
  57. 15.
    Institutional review board (IRB) approvals or equivalent for research with human subjects
  58. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  59. Answer: [N/A]
  60. Justification: Our work does not involve crowdsourcing nor research with human subjects.
  61. 16.
    Declaration of LLM usage
  62. Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research? Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
  63. Answer: [N/A]
  64. Justification: We do not use an LLM to develop any part of the method.

### v3

@@ F.1 Architecture
    Goal buffer $\mathcal{B}$. Stores a diverse set of landmark states from the offline dataset. During planning, the buffer retrieves the nearest stored state $\mathcal{B}(s_{t},z)\rightarrow s_{g}$, to the predicted representation $z$, acting as a discrete decoder. This eliminates hallucination errors by constraining subgoals to states that actually exist in the data.
F.2 Training Objectives
The agent is trained using offline datasets collected by noisy expert policies [park2024ogbench]. All modules are trained jointly using samples from the replay buffer.
Manager Value Function
The manager value function $V^{h}(s_{t},s_{g})$ evaluates the quality of high-level plans. Following our tree-structured return formulation (Sec. 2.3.3), we compute values using the $\min$-operator.
For a given task $(s_{t},s_{g})$ decomposed via a midway state $s_{w}$:
$V^{h}(s_{t},s_{g})=\min(R_{\text{left}}+\gamma^{h}V^{h}(s_{t},s_{w}),R_{\text{right}}+\gamma^{h}V^{h}(s_{w},s_{g}))$
where $R_{\text{left}},R_{\text{right}}$ are binary rewards $\{0,1\}$ indicating whether each subtask is reachable within the threshold of subgoal step (replacing the imagination-based check with a distance threshold).
The value function is trained using an expectile regression objective (similar to IQL [kostrikov2021offline], HIQL [park2023hiql]) to avoid overestimation. However, unlike standard IQL, which regresses toward Bellman backup values, we regress toward our tree-structured return $G^{h}$:
$G^{h}(s_{t},s_{g};s_{w})=\min(R_{\text{left}}+\gamma^{h}\bar{V}^{h}(s_{t},s_{w}),R_{\text{right}}+\gamma^{h}\bar{V}^{h}(s_{w},s_{g}))$
This is the 1-step tree return from Equation 5, where we use target networks $\bar{V}^{h}$ for stability. The loss is then:
$L_{V^{h}}=\mathbb{E}(s_{t},s_{w},s_{g})[\mathcal{L}\tau(G^{h}(s_{t},s_{g};s_{w})-V^{h}(s_{t},s_{g}))]$
where $\mathcal{L}_{\tau}$ is the asymmetric expectile loss:
$\mathcal{L}_{\tau}(u)=|\tau-1[u<0]|\cdot u^{2}$
with expectile $\tau=0.7$. This asymmetric loss causes the value function to approximate an upper expectile of the tree-structured return distribution, avoiding overestimation while maintaining optimism for planning.
Manager Actor
The manager policy learns to predict subgoal representations that maximize the advantage:
$\pi_{M}(z|s_{t},s_{g})\propto\exp(\beta\cdot A^{h}(s_{t},s_{w},s_{g}))$
where $s_{w}$ is the midway state between $(s_{t},s_{g})$ in the dataset trajectory, and the advantage is:
$A^{h}(s_{t},s_{w},s_{g})=\min(V^{h}(s_{t},s_{w}),V^{h}(s_{w},s_{g}))-V^{h}(s_{t},s_{g})$
This encourages the manager to predict subgoals that create balanced, high-value decompositions. The policy is trained via advantage-weighted regression (AWR):
$\mathcal{L}_{\pi_{M}}=-\mathbb{E}_{(s_{t},s_{w},s_{g})}[\exp(\beta\cdot A^{h})\cdot\log\pi_{M}(\phi(s_{t},s_{w})|s_{t},s_{g})]$
where $\beta=3.0$ is the temperature parameter controlling the strength of advantage weighting.
Worker Value Function
The worker value function $V^{l}(s_{t},z)$ where $z=\phi(s_{t},s_{g})$ evaluates how well the agent can reach subgoal $s_{g}$ from state $s_{t}$. It is trained using standard IQL expectile regression:
$\mathcal{L}_{V^{l}}=\mathbb{E}_{(s_{g},s^{\prime},s_{g})}[\mathcal{L}_{\tau}(r+\gamma^{l}V^{l}(s_{t},z)-V^{l}(s_{t},z))]$
where $r=\text{Float}[d(s^{\prime},s_{g})\leq 1]$ is a sparse binary reward indicating goal achievement, and $\gamma^{l}=0.99$ is the discount factor. Similar to HIQL [park2023hiql], the goal representation module $\phi$ is trained using the gradients from the value function loss.
Worker Actor
The worker policy is trained to imitate actions that have high advantages:
$L_{\pi_{W}}=-\mathbb{E}(s,a,s^{\prime},s_{t+K})[\exp(\beta\cdot(V^{l}(s^{\prime},z)-V^{l}(s_{t},z)))\cdot\log\pi_{W}(a_{t}|s_{t},z)]$
where $z=\phi(s_{t},s_{t+K})$. This encourages the worker to execute actions that make progress toward the subgoal $s_{t+K}$.
Goal Buffer Update
The goal buffer is populated and maintained using farthest point sampling (FPS) to ensure diversity:
  1. 1.
    Initialization: Sample $N$ candidate goals from the offline dataset
  2. 2.
    Diversity selection: Every $10000$ steps, we iteratively select observations from the data batch that maximize minimum distance to already-stored states: $\mathcal{B}\leftarrow\arg\max_{s\in\text{data[obs]}}\min_{s^{\prime}\in\mathcal{B}}d(s,s^{\prime})$ where $d$ is a value-based distance estimate $V^{l}(s,s^{\prime})$. This objective selects the set of states that are most temporally distant.
The buffer capacity is set to 2048 goals, which we found to work well for all tasks. But it can be task-dependent. Using value-based FPS ensures the buffer contains goals at varying difficulty levels, not just perceptually diverse states.
F.3 Inference
During inference, the manager recursively decomposes tasks up to depth $D=8$. For each decomposition, the manager predicts a goal representation $\pi_{M}(s_{t},s_{g})\rightarrow z$. Then, the $z$ is used to retrieve the nearest state by comparing the goal representations of the stored states $\phi(s_{t},s_{i})\forall s_{i}\in\mathcal{B}$ with the predicted $z$ using the mean-squared error. The first reachable subgoal in the decomposed subgoal stack (say $s_{w}$) is used as the worker’s subgoal to predict the action as: $\pi_{W}(s_{t},\phi(s_{t},s_{w}))$. Reachability is determined by the normalized value criterion $\tilde{V}^{l}(s_{t},z)>\theta$, where $\tilde{V}^{l}=(V^{l}-\mu)/\sigma$ is the value estimate normalized by running statistics $(\mu,\sigma)$ collected during training for valid subgoals ($s_{t+K}$). We set $\theta=-2.0$, which corresponds to subgoals that are in the top $97.5\%$ of value estimates, preventing over-optimism.
F.4 Evaluation
We evaluate the offline DHP agent on the OGBench navigation benchmark [park2024ogbench], which features significantly larger and more complex environments than the 25-room task:
  • AntMaze-Large/Giant: Large continuous mazes requiring precise locomotion control.
  • HumanoidMaze-Large/Giant: High-dimensional humanoid morphology ($17$ DoF) in large mazes, requiring both locomotion and balance.
Table 4 shows that offline DHP achieves state-of-the-art performance, with particularly strong absolute gains on the HumanoidMaze tasks ($33.8\%$ on large, $71.2\%$ on giant compared to HIQL). Our planning method is highly interpretable, as shown in the subgoal visualization in Fig. 15. For video results, please visit https://sites.google.com/view/dhp-video/home. These are the best results on the OGBench humanoid tasks to the best of our knowledge. These results demonstrate that:
  • Core contributions generalize: The discrete reachability paradigm and tree-structured returns transfer successfully to the offline setting.
  • Architecture modularity: Our planning principles work with both online (CVAE+world model) and offline (goal buffer+representation learning) implementations.
  • Scalability: The method handles high-dimensional morphologies and giant environments where the online version struggled.
Task GCBC GCIVL GCIQL QRL CRL HIQL DHP (Ours)
antmaze-large-
navigate-v0 $24\pm 2$ $16\pm 5$ $34\pm 4$ $75\pm 6$ $83\pm 4$ $91\pm 2$ $\mathbf{93.9\pm 1}$
antmaze-giant-
navigate-v0 $\pm 0 0±0$$\pm 0$ $0\pm 0$ $14\pm 3$ $16\pm 3$ $65\pm 5$ $\mathbf{72.3\pm 5}$
humanoidmaze-large-
navigate-v0 $1\pm 0$ $2\pm 1$ $2\pm 1$ $5\pm 1$ $24\pm 4$ $49\pm 4$ $\mathbf{82.8\pm 5}$
humanoidmaze-giant-
navigate-v0 $\pm 0 0±0$$\pm 0$ $0\pm 0$ $1\pm 0$ $3\pm 2$ $12\pm 4$ $\mathbf{83.2\pm 4}$
Table 4: Performance on OGBench navigation tasks.
((a)) Ant Maze Large
((b)) Ant Maze Giant
Figure 15: Samples of plan visualizations at the OGBench navigation tasks. The green points signify the stored states in the buffer $\mathcal{B}$. The black dot specifies the agent’s current position. The yellow dot labeled ’G’ is the final goal. And the subgoals are represented by blue dots, with their indices in the subgoal stack shown next to each. The agent recursively plans until a reachable subgoal is found, then moves towards it.
F.5 Comparison to Online DHP
The offline variant trades exploration flexibility for robustness:
Advantages:
  • Eliminates subgoal hallucinations via buffer-based retrieval
  • Scales to high-dimensional state spaces (humanoid)
  • No world model training required
  • Handles extremely long horizons (giant mazes)
Tradeoffs:
  • Requires pre-collected offline data
  • Subgoal space limited to buffer contents (though 2048 landmarks prove sufficient)
  • Cannot discover entirely novel subgoals beyond training distribution
Both variants validate that our core insight—replacing distance metrics with discrete reachability and using tree-structured advantages—is the key to effective hierarchical planning, independent of the specific architecture used.
Appendix G Lightsout Puzzle
Lightsout is a complex puzzle where the agent is given a binary 2D array representing the state of lights on/off in a grid of rooms. The objective is to turn off the lights in all rooms; but, toggling the lights in room $(i,j)$ toggles the lights of connected rooms [(i1,j),(i+1,j),(i,j1),(i,j+1)][(i-1,j),(i+1,j),(i,j-1),(i,j+1)]. We measure reachability by iterating over all actions to check if the goal state can be reached in one step (errorless check). The agent is required to plan a path from the initial state to the final all-out state.
To assess the planning capability of DHP, the agent is stripped down to just the planning actor and critic, both of which are implemented as simple dense MLPs. The actor outputs directly in the state space. For a grid $L\times L$, the size of the state space is $2^{L^{2}}$. Thus, the agent must choose the right subgoal state of the possible $2^{L^{2}}$ at each step in tree planning. This is more complex than the 25-room task, where the agent had to essentially pick from 25 rooms at each step. The agent plans up to a depth of $D=5$, and all branches must terminate in success (within a depth of $D=5$) in a single attempt to score $1$; otherwise, it scores 0. Another added complication is that some subgoals can be dead ends and may not lead to the goal. We observe that our agent performs decently (path lengths refer to the lengths of the successful paths). Table 5 shows the final results.
$L$ State space size Success rate Path lengths
22 $16$ $100\%$ $2.42\pm 0.56$
33 $512$ $86.47\%$ $3.12\pm 1.66$
Table 5: Performance at the lights-out puzzle.